The 4 year old with meningistis is to receive ceftriazone 750mg IVPB over 30minutes. The pharmacy provided 750mg in 50ml D5W to be infused IVPB through microdrip infusion system(tubing drop factor 60 gtt/ml) At what rate, in gtt/ml, should the nurse program the IVPB pump.

The order is pencillian 50,000 units. The vial reads peniciliian 500,000units.Add 4.3ml to yeild 5ml. How many ml do you give.
Answer #1

Ans. Receiving ceftriazone = 750 mg IVPB over 30 minutes,

So 1 minutes received = 750/30 = 25 ml/minutes.

Oder 750mg in 50 ml ,so 750/50= 15 ml

Microdrip 60 gtt/ml

Formula= D/A*drop factor

25/15*60 =100 gtt/ml

Penicillin oder 50,000 units

Available 500,000 units

Add. 4.3 ml to yeild 5 ml

Formula = D/A*drop factor

50000/500000*5 =0.5ml